OpenStudy (abymartinez15):

PLEASE HELP ME WILL : "A bicycle manufacturer is studying the reliability of one of its models. The study finds that the probability of a brake defect is 4 percent and the probability of both a brake defect and a chain defect is 1 percent. If the probability of a defect with the brakes or the chain is 6 percent, what is the probability of a chain defect?"

OpenStudy (luffingsails):

Consider drawing a Venn diagram?

OpenStudy (luffingsails):



OpenStudy (abymartinez15):

So would chain defect would be 6?

OpenStudy (luffingsails):

No... The problem says: P(chain defect) OR P(brake defect) = 6 P(chain defect) AND P(brake defect) = 1 P(brake defect) = 6 P(chain defect) = 6 - 4 - 1 = 1

OpenStudy (abymartinez15):

So it will be 1?

OpenStudy (ciarán95):

The Venn Diagram which @luffingsails has drawn above is a really good way of visualising this problem. The key probability formula which we need to use here is: \[P(B~or~C) = P(B) + P(C) - P(B~intersect~C)\] where B is the probability of a brake defect and C is the probability of a chain failing.

OpenStudy (abymartinez15):

okay so that formula what is it for?

OpenStudy (kropot72):

P(B} = 0.04 \[\large P(B\cap C)=0.01\] \[\large P(B\cup C)=P(B)+P(C)-P(B\cap C)\] 0.06 = 0.04 +P(C) = 0.01 P(C) = 0.06 - 0.04 + 0.01 P(C) = 0.03 or 3%

OpenStudy (abymartinez15):

So would it be 3%?

OpenStudy (kropot72):

Yes, the probability of a chain defect is 3%,

OpenStudy (abymartinez15):

Thank you to all three of you guys

OpenStudy (ciarán95):

So, rearranging the above formula, the probability of a chain failing is: \[P(C) = P(B~or~C) - P(B) + P(B~intersect~C)\] I think there's a little mistake in the Venn diagram....it should be:  We know that P(B) = 0.04 overall, P(B or C) = 0.06 and P(B and/intersect C) = 0.01. So, subbing these numbers in, you should get the correct answer!